Multiply. and write the answer in simplest form:
4 1/5 x 10/21 x 9/20

Answer by fractalier(6550)

Make all of the fractions improper and cancel where you can...thus
4 1/5 x 10/21 x 9/20 =
21/5 x 10/21 x 9/20 =
1/5 x 10/1 x 9/20 =
18/20 =
9/10